OpenStudy (anonymous):

complete the square and write the equation in standard form and state whether it is ellipse, hyperbolar or circle: 5x^2 + y^2 = 2y + 1

OpenStudy (anonymous):

its a circle

OpenStudy (anonymous):

can you show me why?

OpenStudy (anonymous):

5x^2+y^2-2y=1 5x^2+y^2-2y+1=1+1 5x^2+(y-1)^2=2 thus a circle

OpenStudy (anonymous):

hope it helpd u

OpenStudy (anonymous):

circle